OpenStudy (amonoconnor):

Find y'' at x=0 of: xy + e^y = e. Any and all help is greatly appreciated!

OpenStudy (anonymous):

oops not clear, too many equal signs sorry

OpenStudy (anonymous):

\[y+xy'+e^yy'=0\] that's better

OpenStudy (anonymous):

i get \[y'=-\frac{y}{x+e^y}\] but it is late so check my algebra

OpenStudy (anonymous):

that's right. for y'

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

then for \(y''\) use the quotient rule

OpenStudy (amonoconnor):

Why does the "e" on the right side of the equals sign go to (0) rather than remaining as e? Wouldn't e^1 stay e^1 just as e^x stays e^x?

OpenStudy (anonymous):

\(e\) is a number

OpenStudy (anonymous):

like \(5\) or \(\pi\)

OpenStudy (amonoconnor):

Why is d/dx of e^x treated so specially then, in calculus?

OpenStudy (anonymous):

\(e^x\) is function, the exponential function but \(e\) is just a number

OpenStudy (anonymous):

like the difference between \[f(x)=2^x\] and \(2\)

OpenStudy (anonymous):

if \[f(x)=2^x\] then \[f'(x)=2^x\ln(2)\] whereas if \[f(x)=0\] a constant then \[f'(x)=0\]

OpenStudy (anonymous):

ooops i meant if \[f(x)=2\] then \[f'(x)=0\]

OpenStudy (amonoconnor):

Gotcha... I'm tired too I guess :p Thanks! I'm working through it again, with this "insight" lols. Give me just one sec.

OpenStudy (amonoconnor):

Is the derivative of the denominator (1-(e^y)*dy/dx) ?

zepdrix (zepdrix):

derivative of the denominator? I'm not sure what that means. Do you mean the denominator of the derivative? :o

OpenStudy (amonoconnor):

derivative of the denominator of y', so g(x) I guess. The (x+e^y) term. Is it (1)?

OpenStudy (amonoconnor):

I'm using the quotient rule to find y'' from y'. :)

zepdrix (zepdrix):

Oh oh oh :O

OpenStudy (amonoconnor):

I got y'' = (x+ e^y - y)/[(x+e^y)^2]. Do I realllyyy want to square that bottom, algebraically? Is that what I should do? :/

zepdrix (zepdrix):

Derivative of the denominator: \(\large\rm =(1+e^yy')\) Is that what you had? Addition, ya?

zepdrix (zepdrix):

Yah your numerator is a lil botched up I think :)

OpenStudy (amonoconnor):

Damn it... X| Okay... starting over...

OpenStudy (amonoconnor):

This is why I said (1) at first, about d/dx(x + e^y)

zepdrix (zepdrix):

I won't ruin any of the fun. But I'll at least remind you of the "set up":\[\large\rm y'=\frac{y}{x+e^y}\] \[\large\rm y''=\frac{\color{royalblue}{(y)'}(x+e^y)-y\color{royalblue}{(x+e^y)'}}{(x+e^y)^2}\]

zepdrix (zepdrix):

For that online-calculator to accurately give you a derivative, you probably need to specify that y is a function of x. So you would input it like this: \(\large\rm \frac{d}{dx}[x+e^{y(x)}]\) At least that's what you have to do on Wolfram Alpha

OpenStudy (amonoconnor):

And the derivative of "e^y" is, officially, (e^y)*y' ? Is that correct?

zepdrix (zepdrix):

Mmmm yes, looks good.

OpenStudy (amonoconnor):

I'm getting: \[y'' = \frac{-2y - \frac{x^2e^y}{x+e^y}}{(x+e^t)^2}\]

zepdrix (zepdrix):

*I think* your y's should have canceled out on top. Hmm lemme check my work again.

zepdrix (zepdrix):

\[\large\rm y''=\frac{\color{royalblue}{(y)'}(x+e^y)-y\color{royalblue}{(x+e^y)'}}{(x+e^y)^2}\] \[\large\rm y''=\frac{y'(x+e^y)-y(1+e^yy')}{(x+e^y)^2}\] \[\large\rm y''=\frac{-y-y-ye^yy'}{(x+e^y)^2}\]Nope ^ I forgot the negative from y' :) hehe. You were right there.

zepdrix (zepdrix):

Plugging in for the other y',\[\large\rm y''=\frac{-2y-ye^y\frac{-y}{x+e^y}}{(x+e^y)^2}\]

zepdrix (zepdrix):

Was that just a small typo in your second derivative maybe? I see an x^2 in the second term, clearly it's a y^2 though, ya?

zepdrix (zepdrix):

\[\large\rm y''=\frac{-2y+\frac{y^2 e^y}{x+e^y}}{(x+e^y)^2}\]

zepdrix (zepdrix):

Oh..... they wanted you to evaluate this at a particular point? Oh goodness, you're making this way hard on yourself then! :O We don't need to do all this extra work.

OpenStudy (amonoconnor):

Yeah, whoops, that was a typo! :( And yep, at x=0...

zepdrix (zepdrix):

Let's first find a corresponding y-coordinate for our x=0,\[\large\rm xy + e^y = e\qquad\to\qquad 0y+e^y=e\]So what's the y-coordinate? :)

OpenStudy (amonoconnor):

e-(e^y) ?

zepdrix (zepdrix):

Hmm you're gettin too fancy. So it simplifies to:\[\large\rm e^y=e^1\]

zepdrix (zepdrix):

(Because the other thing was being multiplied by 0.)

zepdrix (zepdrix):

The bases are the same, so the exponents have to be the same. So \(\large\rm y=1\), y a?

OpenStudy (amonoconnor):

Wait, so...\[xy + e^y = e\] Can be evaluated to find "y" by doing: \[(0)y + (e) = (0)\]

OpenStudy (amonoconnor):

?

zepdrix (zepdrix):

This: \(\large\rm xy + e^y = e\) can be evaluated at x=0, \(\large\rm 0y + e^y = e\) which will allow you to find a coordinate pair, (x,y). Namely, the coordinate pair with x=0, (0,y).

OpenStudy (amonoconnor):

Ah... I see.

OpenStudy (amonoconnor):

So now I plug (0,1) into this?

OpenStudy (amonoconnor):

And that is y'', at the point, right?

zepdrix (zepdrix):

Yes :) you can do that. I was going to say that we could have avoided that really nasty process of quotient rule, and plugging in for the first derivative by using that (0,1) information earlier.

zepdrix (zepdrix):

From here:\[\large\rm y'(x,y)=\frac{-y}{x+e^y}\]Evaluating the first derivative at (0,1) gives us,\[\large\rm y'(0,1)=\frac{-1}{0+e^1}=-e^{-1}\]That's your first derivative at (0,1). Hold onto that information, and back up to this point,\[\large\rm y+y'(x+e^y)=0\]You COULD find your second derivative from here, it might be easier :) You're already pretty deep into your problem with your method, so maybe it's not worth switching at this point. But I show you just in case you're curious.

OpenStudy (amonoconnor):

Thank you:) That would be nicer in the future, for sure! So... I'm getting -3/e^2, but the book says 1/e^2.

zepdrix (zepdrix):

\[\large\rm y'+y''(x+e^y)+y'(1+e^yy')=0\]Now DON'T go through the trouble of trying to solve for y'' just yet. First plug in all the information. x=0 y=1 y'=-e^-1\[\large\rm -e^{-1}+y''(0+e^1)-e^{-1}(1+e^1(-e^{-1}))=0\]Hmm the first derivative wasn't as easy to use as like... y'=5 or something :) lol but it's still not too bad.

zepdrix (zepdrix):

The book says 1/e^2? Hmmm I'm getting -1/e^2...

OpenStudy (amonoconnor):

Me too... Ahhh! Ugg.

zepdrix (zepdrix):

\[\large\rm y''(0,1)=\frac{-2(1)+\frac{(1)^2 e^1}{0+e^1}}{(0+e^1)^2}=\frac{-2+1}{e^2}\]Hmm :d

zepdrix (zepdrix):

Imma check the other way that I did it.

OpenStudy (amonoconnor):

Just checked my math and signs again, and am getting -1/x^2 ... :(

OpenStudy (amonoconnor):

*Sorry: -1/e^2

zepdrix (zepdrix):

Ooo I'm getting 1/e^2 from my other approach :o oh boy

OpenStudy (amonoconnor):

oh god.. did I do it wrong again? :/

OpenStudy (freckles):

hey

OpenStudy (freckles):

the mistake is just leaving off the - in the y'

OpenStudy (amonoconnor):

I'm so frustrated with having these little issues where I can get like, 75% of the way through a problem but then can't get the damn answer! It' not like I don't get the foundations, chain rule, product rule, etc., there're these little situational questions where I just ostensibly don't see how pellet works. God.

OpenStudy (freckles):

\[y'=-\frac{y}{x+e^y} \] somehow that - got dropped

OpenStudy (freckles):

just bring down the - to y''

OpenStudy (amonoconnor):

Okay, will do!

OpenStudy (amonoconnor):

Can you guys help me with another one?

zepdrix (zepdrix):

\[\large\rm y''=\frac{\color{royalblue}{(-y)'}(x+e^y)-(-y)\color{royalblue}{(x+e^y)'}}{(x+e^y)^2}\]Yah I missed the negatives in my second derivative setup :( woops.

zepdrix (zepdrix):

Sure :)

zepdrix (zepdrix):

We won't make the same mistakes this time :D

OpenStudy (freckles):

i do agree it is hard being perfect

OpenStudy (freckles):

i was doing some ugly math stuff the other day and it involved adding 2 and 2 and for some reason I put 3 instead of 4 which made my end result a little off

OpenStudy (freckles):

i know that sounds dumb

OpenStudy (amonoconnor):

I know exactly how you feel... Ugg.

zepdrix (zepdrix):

XD